OpenStudy (anonymous):

Wallis's formula question:

OpenStudy (anonymous):

\[\int\limits_{0}^{pi/2}\cos^4\theta d \theta\]

OpenStudy (anonymous):

i have the formula infront of me, but i am alittle confused as how to apply it, in my book they applied the formula and found this integral to be :3pi/16

OpenStudy (anonymous):

oh i just figured it out, wallis formula goes like this: If n is even (n greater than or equal to 2), then: \[\int\limits_{0}^{pi/2}\cos^n \theta d \theta\]=(1/2)(3/4)(5/6).....((n-1)/(n))(pi/2)

OpenStudy (anonymous):

now 1/2 corrspond to n=2, and then from then (1/2) time(3/4) corresponds to n=4, (since n has to be even)

OpenStudy (anonymous):

thus for my integral i have (1/2)(3/4)(pi/2)=3theta/8

OpenStudy (anonymous):

nice and easy way to integrate these sin and cos functions with even powers

OpenStudy (beginnersmind):

Nice, haven't seen it before. Reading the wikipedia article now. Seems like it's more important theoretically, as the integral itself isn't too hard.

OpenStudy (anonymous):

yeah, sure, but when your taking a test and your time is running low, its good to have this in your arsenal

OpenStudy (beginnersmind):

Hm, I guess it could come up often when you're trying to calculate Fourier series.